OpenStudy (anonymous):

For each set of numbers, determine whether the numbers represent the lengths of the sides of an acute triangle, a right triangle, an obtuse triangle, or no triangle. a. √32 √26 √8 b. 6, 8, 10 c. 7.1, 10.8, 23.3

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (cwrw238):

b) is a right angled triangle because 10^2 = 6^2 + 8^2 a) squaring each side = 32 . 26 and 8 32 is less than 26 + 8 so this is acute angled triangle c) this is not a triangle because 7.1 + 10.8 < 23.3